1. A radioactive substance plutonium 239 has a half life of 24100 years. Initially, there is 30mg of the substance, find how much will remain for the first 1000 years. How long for the substance to decay 90% of its initial mass? Ans. 29.15mg; 80,058.47 years
  2.  A thermometer reading of 50°C was plunged into a tub of frozen water. If the thermometer reads 30° after 5 seconds, what will be the reading after 10 seconds? How long will it take for the thermometer reading to drop to 20°C? Ans. 18°C; 8.97 seconds
1
Expert's answer
2021-11-12T16:08:20-0500

1.


"R=R_0e^{kt}"

A half life is 24100 years


"\dfrac{1}{2}R_0=R_0e^{k(24100)}"

"k(24100)=-\ln 2"

"k=-\ln 2/24100"

"R=R_0(2)^{-t\/24100}"

Initially "R_0=30 mg".

(i)


"R(1000)=300(2)^{-1000\/24100}"

"R(1000)=29.15mg"

(ii)


"R=R_0(2)^{-t\/24100}=0.1R_0"

"\ln((2)^{-t/24100})=\ln(0.1)"

"-t/24100=\ln (0.1)/\ln(2)"

"t=24100(\ln(10)/\ln(2))"

"t=80058.467\ years"

"t=80058.47\ years"

2. The temperature of a body at any time "t" is


"T_b=(T_{b0}-T_m)e^{kt}+T_m"

Initially, ܶ"T_{b0}=50\degree C" and "T_m=0\degree C"

Given "T_b(5)=30\degree C"


"30=(50-0)e^{5k}+0"

"e^{5k}=0.6"

(i)


"T_b(10)=(50-0)e^{10k}+0=50(0.6)^2"

"T_b(10)=18\degree C"

(ii)

"k=0.2\ln(0.6)"

"T_b(t_1)=(50-0)e^{0.2\ln(0.6)t_1}+0=20"

"e^{0.2\ln(0.6)t_1}=0.4"

"t_1=\dfrac{5\ln(0.4)}{\ln(0.6)}"

"t_1=8.97\ sec"